If p is directly proportional to q, and p=18 when q=3, find q when p is 48
Answer by John10(297)

because p is directly proportions to q then the formula is:
p = k * q where k is the variance constant
If p = 18 and q = 3 then
k = p/q = 18/3 = 6
Then we find the new q:
p = 6 * q then q = p/6 = 48/6 = 8
John10:)